What is the strength of the ionic interactions between an ion with 2 positive charges and an ion with 1 negative charges, when the distance between them is 15.5 nanometers and the solvent is benzene, with a dielectric constant of 4.6?

1
Expert's answer
2022-03-07T11:45:11-0500

According to the Coulomb's law, the force is given as follows:


"F = k\dfrac{|q_1q_2|}{\varepsilon r^2}"

where "k = 9\times 10^9N\cdot m^2/C^2" is the Coulomb's constant, "q_1 = +2e" is the charge of the first ion, "q_2 = -e" is the charge of the second ion (here "e = 1.6\times 10^{-19}C" is the elementary charge), "r = 15.5\times 10^{-9}m" is the distance between ions, and "\varepsilon = 4.6" is the dielectric constant of benzen.

Thus, obtain:


"F = k\dfrac{2e^2}{\varepsilon r^2} = 9\cdot 10^9\cdot \dfrac{2\cdot (1.6\times 10^{-19})^2}{4.6\cdot (15.5\times 10^{-9})^2} \approx 4.2\times 10^{-13}N"

Answer. "4.2\times 10^{-13}N".